3) Convert following decimal to 8-bit signed numbers in hexadecimal, use two’s-complement for signed integer

127d, -20d, -128d, -1d

4) Convert the 16-bit signed numbers to the decimal

C0A3h, 3AECh, 0101 1001 0111b, 1011 0101 1001 0111b

please solve the problems step by step. It would be of great help.
Answer #1

If you need any corrections or clarifications kindly comment

Please give a Thumps Up if you like the answer.

3) Convert following decimal to 8-bit signed numbers in hexadecimal

Steps for converting decimal to hexadecimal number(unsigned numbers):

  1. Start dividing the decimal number by 16.

  2. Note down the remainder in decimal and in hexadecimal.

  3. Again divide the quotient by 16.

  4. Repeat step 2 and 3 until quotient is equal to 0.

  5. The hexadecimal value is the sequence of the remainders in hex from the last to first.

HEXADECIMAL

0

1

2

3

4

5

6

7

8

9

A

B

C

D

E

F

DECIMAL

0

1

2

3

4

5

6

7

8

9

10

11

12

13

14

15

a) 127 d=7F in hexadecimal

DIVISION

RESULT

REMAINDER

REMAINDER
(in HEX)

127 / 16

7

15

F

7/ 16

0

7

7

ANSWER

7F

Steps for converting signed decimal to hexadecimal

  1. First convert the decimal number to two's complement signed binary.
  2. Then convert that binary value to hexadecimal.

b) -20d

Step1: Convert 2010 to binary:

Division
by 2

Quotient

Remainder

Bit #

20/2

10

0

0

10/2

5

0

1

5/2

2

1

2

2/2

1

0

3

1/2

0

1

4

So 2010 = 0101002

Fill the value to 16 bits by adding 10 leading zeroes on the left

Thus 2010 = 0000 0000 0001 01002

Step2: Convert to 1’s complement by inverting the bits

Result=1111 1111 1110 1011

Step 3: Find 2’s complement by adding 1 to the above result

Result=1111 1111 1110 1011+1=1111 1111 1110 1100

Step 4: Convert this binary value to hexadecimal. For that group the bits into groups of four, then convert each group to its hexadecimal equivalent

1111 1111 1110 1100b=FFEC in hexadecimal

b) -128d

Step1: Convert 12810 to binary:

Division
by 2

Quotient

Remainder

Bit #

128/2

64

0

0

64/2

32

0

1

32/2

16

0

2

16/2

8

0

3

8/2

4

0

4

4/2

2

0

5

2/2

1

0

6

1/2

0

1

7

So 12810 = 1000 00002

Fill the value to 16 bits by adding 8 leading zeroes on the left.

Thus 12810 = 0000 0000 1000 00002

Step2: Convert to 1’s complement by inverting the bits

Result=1111 1111 0111 1111

Step 3: Find 2’s complement by adding 1 to the above result

Result=1111 1111 0111 1111+1=1111 1111 1000 0000

Step 4: Convert this binary value to hexadecimal. For that group the bits into groups of four, then convert each group to its hexadecimal equivalent

1111 1111 1000 0000=FF80 in hexadecimal

d) -1d

Step1: Convert 110 to binary:

Division
by 2

Quotient

Remainder

Bit #

1/2

0

1

0

So 110 = 12

Fill the value to 16 bits by adding 15 leading zeroes on the left

Thus 110 = 0000 0000 0000 00012

Step2: Convert to 1’s complement by inverting the bits

Result=1111 1111 1111 1110

Step 3: Find 2’s complement by adding 1 to the above result

Result=1111 1111 1111 1110+1=1111 1111 1111 1111

Step 4: Convert this binary value to hexadecimal. For that group the bits into groups of four, then convert each group to its hexadecimal equivalent

1111 1111 1111 1111b=FFFF in hexadecimal

4) Convert the 16-bit signed numbers to the decimal

a. C0A3h=49315d

Ch=12d

Ah=10d

C0A3=12*163+0*162+10*161+3*160

        =12*4096+0+10*16+3*1

       =49152+0+160+3

       =49315

b. 3AECh=15084d

Ch=12d

Eh=14d

Ah=10d

3AEC=3*163+10*162+14*161+12*160

        =3*4096+10*256+14*16+12*1

       =12288+2560+224+12

       =15084

c.0101 1001 0111b=1431d

0101 1001 0111=(0*211)+(1*210)+(0*29)+(1*28)+(1*27)+(0*26)+(0*25)+(1*24)+(0*23)+(1*22)+(1*21)+(1*20)

=0+1024+0+256+128+0+0+16+0+4+2+1

=1431

d.

1011 0101 1001 0111b=46487d

1011 0101 1001 0111b=

(1*215)+(0*214)+(1*213)+(1*212)+(0*211)+(1*210)+(0*29)+(1*28)+(1*27)+(0*26)+(0*25)+(1*24)+(0*23)+(1*22)+(1*21)+(1*20)

=32768+0+8192+4096+0+1024+0+256+128+0+0+16+0+4+2+1

=46487